Can someone help me solve this problem?
I need to multiply:
(-3)(-6) + (4)(-2)
Thanks,
Lyn

Answer by fractalier(2101)

Here we have
(-3)(-6) + (4)(-2) =
18 + (-8) =
18 - 8 =
10